Head-First Company plans to sell 5,800 bicycle helmets at $67 each in the coming year. Variable cost is 54% of the sales price; the contribution margin is 46% of the sales price. Total fixed cost equals $56,235 (includes fixed factory overhead and fixed selling and administrative expense).

Required: 1. Calculate the sales revenue that Head-First must make to break even by using the break-even point in the sales equation.

2. Check your answer by preparing a contribution margin income statement based on the break-even point in sales dollars.
Answer #1

1.

Given that comtirbtuion margin is 46% of sales

Ie, Contribution = Sales*46%

At Break Even, Contribution = Fixed Cost

So, 56,235 = Sales * 46%

Hence, Sales = 56,235/46% = $ 122,250.

Break Even (Units to be sold) = 122,250/67 => 1824.63 or 1825 units.

2.

Contribution Format Income Statement ($)
Sales 122,250.00
Less: Variable Cost (122,250*54%) 66,015.00
Contribution 56,235.00
Less: Fixed cost 56,235.00
Income/(Loss) 0.00